# LayerDAG: A Layerwise Autoregressive Diffusion Model for Directed Acyclic Graph Generation

### v1

@@ 3.3 Permutation invariance and model generalization
Both bidirectional MPNN layers and positional encodings of layer depth are permutation equivariant, and sum/mean pooling operators are permutation invariant. Then, it follows that for any permutation $\Pi$, we have $\mathbb{P}\left(\Pi^{(l+1)}(G^{(l+1)})\mid\Pi^{(\leq l)}(G^{(\leq l)})\right)=\mathbb{P}\left(G^{(l+1)}\mid G^{(\leq l)}\right)$, where $\Pi^{(l+1)}$ corresponds to $\mathcal{V}^{(l+1)}$ and $\Pi^{(\leq l)}$ corresponds to $\mathcal{V}^{(\leq l)}$. Finally, $\mathbb{P}\left(G\right)=\prod_{l=0}^{L-1}\mathbb{P}\left(G^{(l+1)}\mid G^{(\leq l)}\right)$ is also permutation invariant.
@@ 3.4 Flexible quality-efficiency trade-off with a layer-index-based denoising schedule
Note that as $l$ increases, both $|\mathcal{V}^{(\leq l)}|$ and $|\mathcal{E}^{(l+1)}|$ increase in general, resulting in more complex edge dependencies. To effectively handle this pattern, we introduce a non-uniform denoising schedule for better allocation of the time budget. Specifically, we propose to set the total number of denoising steps to linearly increase in $l$.
@@ 5 Experiments
Baselines We consider three non-diffusion autoregressive baselines – GraphRNN, D-VAE, and GraphPNAS. GraphRNN originally tackles undirected graph structure generation (You et al., 2018), and we adopt an extension of it for attributed DAG generation (Zhang et al., 2019). D-VAE is a variational auto-encoder and employs a nodewise autoregressive decoder. Both GraphRNN and D-VAE employ topological orderings and one-hot encoding of node IDs (Zhang et al., 2019). GraphPNAS sequentially generates incident edges for a node set of constant size, using a mixture of Bernoulli distributions to model intra-set dependencies. To compare the capability of LayerDAG in set generation against GraphPNAS, we further adapt GraphPNAS by using mixtures of multinoulli distributions for generating a set of node attributes and setting the node set size to be the averaged layer size. For pure diffusion baselines, we implement OneShotDAG, a non-autoregressive variant of LayerDAG. For ablation study of multiple rounds of refinement, we report results with a single denoising step, denoted by LayerDAG ($\mathbf{T=1}$). For details of model extensions, see Appendix B.
@@ 5.1 Generating synthetic DAGs with strong logical rules (Q1)
We use LP with different $\rho$’s to generate the datasets $D_{\rho}$ and train different generative models based on $D_{\rho}$. After training, we use each model to generate DAGs of the same number as that in $D_{\rho}$. To evaluate the hard logical constraints, we assess the validity of generated DAGs by measuring the proportion of generated DAGs that satisfy the imposed hard constraints. To evaluate the soft constraints, we compare the distributions of graph statistics between generated DAGs and an equal number of real DAGs. We measure the $1$-Wasserstein distance ($W_{1}$) between the distributions of layer numbers ($L$) for the two graph sets. Additionally, we measure Maximum Mean Discrepancy (MMD) (You et al., 2018) between two sets of graph statistic distributions corresponding to the graph sets. Specifically, we report MMD for the distributions of layer size ($|\mathcal{V}^{(l)}|$).
@@ 5.2 Conditional generation for real-world datasets from diverse computational platforms (Q2)
Datasets. We repurpose and adapt three representative real-world datasets, originally developed for DAG property prediction, to serve as testbeds for conditional DAG generation. The datasets are associated with computation workloads executed on diverse hardware platforms, and they well fit the end scenario of synthetic data sharing for system/hardware benchmarking. Originally released as part of the TpuGraphs dataset, TPU Tile is a collection of kernel graphs for machine learning workload on Tensor Processing Units (TPUs), with graph labels $y$ indicating the runtime averaged over a set of compilation configurations (Phothilimthana et al., 2023b). High-level synthesis (HLS) is a collection of data flow intermediate representation graphs for compiled C programs, with each DAG labeled according to the resource usage of look up table measured on Field Programmable Gate Arrays (FPGAs) (Wu et al., 2022). NA-Edge is a collection of DAGs representing neural architectures, with labels indicating their inference latency on mobile CPU (Dong & Yang, 2020; Zhang et al., 2021). We follow Wang et al. (2024) for data pre-processing. Table 2 presents the dataset statistics. We perform a random train/val/test split for all datasets. For more details on dataset adaptation and data statistic distributions, see Appendix E.
@@
Dataset # graphs $|\mathcal{V}|$ $\max|\mathcal{V}|$ $|\mathcal{E}|$ $\max|\mathcal{E}|$ $L$ $\max L$ $|\mathcal{V}^{(l)}|$ $\max|\mathcal{V}^{(l)}|$ attributes label info TPU Tile 6, 301 40.8 394 42.9 711 11.2 72 3.6 21 1 TPU runtime HLS 2, 062 88.6 356 110.7 477 27.8 78 3.2 28 7 FPGA resource usage NA-Edge 2, 000 231.1 339 265.8 385 149.1 185 1.5 4 14 mobile CPU inference latency
@@ 5.3 Label generalization in conditional generation (Q3)
To assess the model capabilities in generalizing to unseen regimes of label values, we establish a challenging out-of-distribution testbed using the TPU Tile dataset. We sort the real-world DAGs based on their labels and then partition them into five quantiles. For an experiment, we exclude one quantile of data during the development of the generative models. After training the generative models, we conduct a random train/val/test split on the previously excluded quantile. We evaluate conditional generation exclusively on this quantile, i.e., generating DAGs conditioning on labels from this quantile. Essentially, this setup requires the models to perform interpolation for intermediate quantiles and extrapolation for the end quantiles. An analysis of the TPU tile dataset reveals a long-tailed label distribution, with larger value ranges in latter quantiles, detailed in Appendix E. Therefore, we choose the 5th quantile for an extrapolation study and the 4th quantile for an interpolation study, which presents significant challenges. Besides the previously employed BiMPNN model, we leverage another surrogate model that stood out as a top-5 solution from more than 600 submissions in a Kaggle competition for the TpuGraphs dataset (Phothilimthana et al., 2023a; Kag, 2023).
Table 4 presents the evaluation results. As expected, the extrapolation setting is extremely challenging, where most models even failed to achieve a single positive Pearson correlation coefficient. In contrast, LayerDAG consistently achieves the best metrics for ML-based evaluation in this setting, with a gain of at most about $0.2$ in Pearson correlation coefficient for BiMPNN. Overall, LayerDAG significantly outperforms the rest models for both the extrapolation and interpolation settings, suggesting a superior generalization capability arising from the combination of autoregressive layerwise generation and diffusion.

### v2

@@ 3.3 Permutation invariance and model generalization
Bidirectional MPNN layers are permutation equivariant, and sum and mean pooling operators are permutation invariant. It follows that for any permutation $\Pi$, we have $\mathbb{P}\left(\Pi^{(l+1)}(G^{(l+1)})\mid\Pi^{(\leq l)}(G^{(\leq l)})\right)=\mathbb{P}\left(G^{(l+1)}\mid G^{(\leq l)}\right)$, where $\Pi^{(l+1)}$ corresponds to $\mathcal{V}^{(l+1)}$ and $\Pi^{(\leq l)}$ corresponds to $\mathcal{V}^{(\leq l)}$. Finally, $\mathbb{P}\left(G\right)=\prod_{l=0}^{L-1}\mathbb{P}\left(G^{(l+1)}\mid G^{(\leq l)}\right)$ is also permutation invariant.
@@ 3.4 Flexible quality-efficiency trade-off with a layer-index-based denoising schedule
Note that as $l$ increases, both $|\mathcal{V}^{(\leq l)}|$ and $|\mathcal{E}^{(l+1)}|$ increase in general, resulting in more complex edge dependencies. To effectively handle this pattern, we introduce a non-uniform denoising schedule for better allocation of the time budget. Specifically, we propose to set the total number of denoising steps for a layer to linearly increase in $l$.
@@
The key assumption is that more complex patterns emerge as the layer depth increases, such as long range dependencies between a pair of layers with a big difference in layer depth, which calls for the use of more denoising steps. When this does not hold, as in the scenario of trees with stationary layer distributions, this layer-based linear denoising schedule will not be useful.
@@ 5 Experiments
Baselines We consider three non-diffusion autoregressive baselines – GraphRNN, D-VAE, and GraphPNAS. GraphRNN originally tackles undirected graph structure generation You et al. (2018), and we adopt an extension of it for attributed DAG generation Zhang et al. (2019). D-VAE is a variational auto-encoder and employs a nodewise autoregressive decoder. Both GraphRNN and D-VAE employ topological orderings and one-hot encoding of node IDs Zhang et al. (2019). GraphPNAS sequentially generates incident edges for a node set of constant size, using a mixture of Bernoulli distributions to model intra-set dependencies. To compare the capability of LayerDAG in set generation against GraphPNAS, we further adapt GraphPNAS by using mixtures of multinoulli distributions for generating a set of node attributes and setting the node set size to be the averaged layer size. For pure diffusion baselines, we implement OneShotDAG, a non-autoregressive variant of LayerDAG. For the ablation study of multiple rounds of refinement, we report results with a single denoising step, denoted by LayerDAG ($\mathbf{T=1}$). For details of model extensions, see Appendix B.
@@ 5.1 Generating synthetic DAGs with strong logical rules (Q1)
We use LP with different $\rho$’s to generate the datasets $D_{\rho}$ and train different generative models based on $D_{\rho}$. After training, we use each model to generate DAGs of the same number as that in $D_{\rho}$. To evaluate the hard logical constraints, we assess the validity of generated DAGs by measuring the proportion of generated DAGs that satisfy the imposed hard constraints. To evaluate the soft constraints, we compare the distributions of graph statistics between generated DAGs and an equal number of real DAGs. We measure the $1$-Wasserstein distance ($W_{1}$) between the distributions of layer numbers ($L$) for the two graph sets. Additionally, we measure Maximum Mean Discrepancy (MMD) You et al. (2018) between two sets of graph statistic distributions corresponding to the graph sets. Specifically, we report MMD for the distributions of layer size ($|\mathcal{V}^{(l)}|$).
@@ 5.2 Conditional generation for real-world datasets from diverse computational platforms (Q2)
Datasets. We repurpose three representative real-world datasets, originally developed for DAG property prediction, to serve as testbeds for conditional DAG generation. The datasets are associated with computation workloads executed on diverse hardware platforms, and they well fit the end scenario of synthetic data sharing for system/hardware benchmarking. Originally released as part of the TpuGraphs dataset, TPU Tile is a collection of kernel graphs for machine learning workload on Tensor Processing Units (TPUs), with graph labels $y$ indicating the runtime averaged over a set of compilation configurations (Phothilimthana et al., 2023b). High-level synthesis (HLS) is a collection of data flow intermediate representation graphs for compiled C programs, with each DAG labeled according to the resource usage of look up table measured on Field Programmable Gate Arrays (FPGAs) (Wu et al., 2022). NA-Edge is a collection of DAGs representing neural architectures, with labels indicating their inference latency on mobile CPU (Dong & Yang, 2020; Zhang et al., 2021). Table 2 presents the dataset statistics. We perform a random train/val/test split for all datasets. For more details on dataset adaptation and data statistic distributions, see Appendix E.
@@
Dataset # graphs $|\mathcal{V}|$ $\max|\mathcal{V}|$ $|\mathcal{E}|$ $\max|\mathcal{E}|$ $L$ $\max L$ $|\mathcal{V}^{(l)}|$ $\max|\mathcal{V}^{(l)}|$ attributes label info TPU Tile 6, 301 40.8 394 42.9 711 11.2 72 3.6 21 1 TPU runtime HLS 2, 062 88.6 356 110.7 477 27.75 78 3.2 28 7 FPGA resource usage NA-Edge 2, 000 231.1 339 265.8 385 149.1 185 1.5 4 14 mobile CPU inference latency
@@ 5.3 Label generalization in conditional generation (Q3)
To assess the model generalization to unseen regimes of label values, we establish a challenging out-of-distribution testbed using the TPU Tile dataset. We sort the real-world DAGs based on their labels and then partition them into five quantiles. For an experiment, we exclude one quantile of data during the development of the generative models. After training the generative models, we conduct a random train/val/test split on the previously excluded quantile. We evaluate conditional generation exclusively on this quantile, i.e., generating DAGs conditioning on labels from this quantile. Essentially, this setup requires the models to perform interpolation for intermediate quantiles and extrapolation for the end quantiles. An analysis of the TPU tile dataset reveals a long-tailed label distribution, with larger value ranges in latter quantiles, detailed in Appendix E. Therefore, we choose the 5th quantile for an extrapolation study and the 4th quantile for an interpolation study, which presents significant challenges. Besides the previously employed BiMPNN model, we leverage another surrogate model that stood out as a top-5 solution from more than 600 submissions in a Kaggle competition for the TpuGraphs dataset (Phothilimthana et al., 2023a; Kag, 2023).
Table 4 presents the evaluation results. As expected, the extrapolation setting is particularly difficult, with most baselines failing to attain a single positive Pearson correlation. By contrast, LayerDAG consistently outperforms competitors, improving BiMPNN’s correlation by up to 0.2. While 0.2 remains modest for practical usage, the combination of autoregressive layerwise generation and diffusion drives LayerDAG ’s stronger generalization across both settings.